Proposal:
K Kids
Target of Oportunity:
No
Time Observed:
8.75
Time Engineering:
0.0
Time lost technical:
0.25
Time lost weather:
0.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
9

- 00:42 UT Fails to start exposure on T0167661160, rebooted OPTGUI. OK? No, Images 119 and 1120 unbinned PLUS system changed timestamps to 21:05 and 21:12 local time! After a new reboot, a third 240 s image is finally ok. Guess T0167661160 needs to be rescheduled.
- DIMM seeing reached a 0.25" value (@ 04:33UT)... or the instrument is joking.
- Twice tonight (TOI-1015 and TIC40079924) the telescope fails to reach position, apparently erring in Declination. TCS reports OK but clearly is not, so it is needed to send it again. Encoder?
- After 08:15 UT, thin clouds moving from south west to N, NE, likely impacts only chi Cen and no previous object.
- NOS coordinates of HR 5285 = chi Cen are wrong, used those at CDS.
- No time for 6233126130764332416. Completed last integration (#6 out of requested 12) on chi Cen with the Sun at -9.1 degrees of elevation.